A point charge, -30µC, is placed at the origin of coordinates. Find the electric field at the point x=5m on the x-axis
1
Expert's answer
2018-03-28T12:11:07-0400
An electric field at the point x=5m: E=kq/x^2 =(9*〖10〗^9*30*〖10〗^(-6))/5^2 =10.8*〖10〗^3 N/C